OpenStudy (anonymous):

if 9,8 is a endpoint of a segment and 4,12 is the midpoint, find the other endpoint

OpenStudy (anonymous):

-1,16

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

if we know a distance and direction to go this would be eaiser; finish - start = distance and direction we start at a point and end at the mid

OpenStudy (anonymous):

\[(x _{2}-x _{1})/2=4=>9-x _{1}=8=>x _{1}=1\] Do the same for y.

OpenStudy (amistre64):

m-e = distance and direction from end to mid m+(m-e) = the rest of the way

OpenStudy (amistre64):

m = 4,12 2m = 8,24 e = 9,8 -e = -9, -8 2m = 8, 24 -e = -9, -8 ----------- e = -1,16

OpenStudy (anonymous):

Let the coordinates of the first endpoint (x1,y1) and the coordinates of the other be (x2,y2). Then ther coordinates of the midpoint is (x1+x2/2,y1+y2/2)

OpenStudy (anonymous):

Nw if we take the coordinates of the other endpoint as (x,y) 9+x/2=4 9+x=8 x=-1 8+y/2=12 8+y=24 y=16

OpenStudy (anonymous):

Is that clear nw?

OpenStudy (anonymous):

madskills?

OpenStudy (anonymous):

awsome, thanks alot!!

OpenStudy (anonymous):

yw